OpenStudy (anonymous):

Find the limit as x approaches -3 of (sqrt(6-x) - 3)/(x+3)

OpenStudy (anonymous):

I know you multiply by the conjugate, but when I did that nothing factored out.

OpenStudy (anonymous):

Oh, then I don't know. @Hero Please help.

hero (hero):

Try multiplying by \(\dfrac{\sqrt{6 - x} + 3}{\sqrt{6 - x } + 3}\)